OpenStudy (anonymous):

PLEASE..Help me figure out if there is ….. A. If there is one solution? (type and ordered pair) B. If there are infinitely many solutions? C. Or if there are no solutions. Use the method of substitution to solve the system of linear equations. 6a=7-b 6a=4-b

OpenStudy (lgbasallote):

subtract the two equations...what do you get?

OpenStudy (anonymous):

i am really new to this so... 6a=3b?

OpenStudy (anonymous):

i am unsure of the steps.

OpenStudy (lgbasallote):

6a=7-b 6a=4-b what is 6a - 6a?

OpenStudy (anonymous):

0?

OpenStudy (lgbasallote):

right. and what's 7 - 4?

OpenStudy (anonymous):

3

OpenStudy (lgbasallote):

and -b - (-b)?

OpenStudy (anonymous):

0?

OpenStudy (lgbasallote):

yes

OpenStudy (lgbasallote):

so.. 6a=7-b 6a=4-b ===== 0 = 3 + 0 is this true? is 0 = 3+0?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

what is 3 + 0?

OpenStudy (anonymous):

3

OpenStudy (lgbasallote):

so is 0 =3??

OpenStudy (anonymous):

so that would be false right?

OpenStudy (lgbasallote):

yes!!

OpenStudy (lgbasallote):

that means there are no solutions

OpenStudy (anonymous):

Thank you!!!!!! I like when people do step by step and help me understand! Instead of giving the answers. I learn better for tests :)

OpenStudy (lgbasallote):

welcome :D and i like people who would rather listen to steps rather than the answers

OpenStudy (anonymous):

:)